OpenStudy (anonymous):

Use the Binomial Theorem complete the binomial expansion. (3x + 5)3 = 27x3 + 135x2 + __________

alones (alones):

is that all in one O.O

OpenStudy (anonymous):

What do you mean? Yeah, it's one binomial. I need to complete it but I'm confused.

OpenStudy (anonymous):

(a + b)^3 = a^3 + b^3 + 3*a*b*(a + b)

OpenStudy (anonymous):

can you solve it?

OpenStudy (anonymous):

I'm still confused, sorry. What do I insert in (a) and (b)? @stefancipe

OpenStudy (anonymous):

you have (3x+5)^3 , so a = 3x, and b= 5

OpenStudy (anonymous):

ah okay i'll try it out

OpenStudy (anonymous):

is it 225x?

OpenStudy (anonymous):

i already solved it, so post solution here so i can check it :)

OpenStudy (anonymous):

yes, but you're missing one small part :) 5^3

OpenStudy (anonymous):

oh, okay!!! i get it, i get it. thank you so much! :)

OpenStudy (anonymous):

no problem :)